Could you please help me with this equation?
x+2y=5
3x-15+-6y
It has to be solved using elimination or substitution.
I need serious help!!!!

Found 2 solutions by stanbon, jcmtnez:
Answer by stanbon(75887)

x+2y=5
3x-15=-6y
Solve the 1st for "x", as follows:
x=5-2y
Substitute that into #2 to get;
3(5-2y)-15=-6y
15-6y-15=-6y
0=0
This says the two equations are really the
same equation.
So, what is the solution?
It is the infinite set of points that constitute the line.
Cheers,
Stan H.

Answer by jcmtnez(53)

I think that you mean:
x+2y=5
3x-15=6y
If that is the system I will show you how to solve it.
By Elimination

(1)x+2y=5 We place the x and the y where they must be
3x-6y=15
(2) We multiply the first equation by 3 in both sides to obtain a similar equation.
3(x+2y)=3(5)
3x-6y=15
(3) 3x+6y=15 Now we obtain this system.
3x-6y=15
(4) Finally we sum both equation obtaining a new one wich only one variable.
3x+6y=15 We obtain 6x=30, then x=5, and therefore y=0
+3x-6y=15

By Substitution
x+2y=5 We leave one variable alone, x=5-2y
3x-6y=15 We substitute the variable in the seconde equation
3(5-2y)-6y=15
15-6y-6y=15
-12y=0 and therefore y=0
Now from the first equation

x+2(0)=5
x+0=5 And therefore x=5

Here are the solutions, I hope that they are useful for you.